OpenStudy (carlyleukhardt):

What is the sum? –19 + 41

OpenStudy (khalilforthewin):

-.- 23

OpenStudy (carlyleukhardt):

Okay i was just making sure.

OpenStudy (carlyleukhardt):

XD

OpenStudy (anonymous):

lol its 22

OpenStudy (carlyleukhardt):

>,<

OpenStudy (anonymous):

its simple math 41-19

OpenStudy (carlyleukhardt):

can u help me with another one @JColon25

OpenStudy (zzr0ck3r):

-19+41 = 41-19

OpenStudy (carlyleukhardt):

What is the value of u + v + w when u = –10, v = 4, and w = 8?